### v1

# SAFES: Sequential Privacy and Fairness Enhancing Data Synthesis for Responsible AI

Abstract
As data-driven and AI-based decision making gains widespread adoption in most disciplines, it is crucial that both data privacy and decision fairness are appropriately addressed. While differential privacy (DP) provides a robust framework for guaranteeing privacy and several widely accepted methods have been proposed for improving fairness, the vast majority of existing literature treats the two concerns independently. For methods that do consider privacy and fairness simultaneously, they often only apply to a specific machine learning task, limiting their generalizability. In response, we introduce SAFES, a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure that sequentially combines DP data synthesis with a fairness-aware data transformation. SAFES allows full control over the privacy-fairness-utility trade-off via tunable privacy and fairness parameters. We illustrate SAFES by combining AIM, a graphical model-based DP data synthesizer, with a popular fairness-aware data pre-processing transformation. Empirical evaluations on the Adult and COMPAS datasets demonstrate that for reasonable privacy loss, SAFES-generated synthetic data achieve significantly improved fairness metrics with relatively low utility loss.
keywords: Differential privacy, Machine learning fairness, Synthetic data
1 Introduction
Data-driven and AI-based decision making are being adopted in many disciplines and data collected as part of this process often contain sensitive data from individuals. These data are frequently used to make socially impactful decisions including, but not limited to, determining who gets approved for a loan, predicting which applicants should be hired for a job, or forecasting which previously convicted criminals will re-offend. While such applications can and do have legitimate benefits, it is of paramount importance to ensure the use of such sensitive data is responsible and carried out with the highest possible ethical standards.
There are two important ethical concerns when working with sensitive personal data in training machine learning (ML) and AI algorithms: privacy and fairness. Even anonymized datasets can be leveraged by attackers to infer masked or removed data (Narayanan and Shmatikov,, 2008; Ahn,, 2015; Sweeney,, 2015) and blackbox access to a model is sufficient to infer membership in the training data (Shokri et al.,, 2017), which can itself be a privacy violation. Knowing that an individual belongs to a dataset used for recidivism prediction, for example, is equivalent to knowing that person was convicted of a crime. The perpetuation of social discrimination, bias, and other forms of unfairness in the decisions made by ML/AI models raises another major ethical concern. One famous previous work, for example, demonstrated the presence of discrimination against darker skin colors in commercial gender classification systems (Buolamwini and Gebru,, 2018). The naïve approach of simply removing the group indicator when training these models has been shown insufficient to properly ensure fairness (Calders and Žliobaitė,, 2013). Fairness has increasingly been recognized in the ML community as a complicated notion and a lot of research has been devoted to defining and ensuring fairness (Verma and Rubin,, 2018).
Since datasets with privacy concerns are likely to also have fairness concerns, and vice versa, it is critical to develop efficient privacy- and fairness-enhancing methods for releasing and analyzing data. In this work, we propose a framework for synthesizing data that simultaneously addresses privacy and fairness concerns by strategically and sequentially combining privacy-preserving data synthesis and fairness-aware pre-processing transformations. We aim to provide a generalized solution that safeguards sensitive personal information, upholds fairness, and keeps the utility of released data close to the original.
There remains a scarcity of research on general-use synthetic data that both satisfies DP guarantees and reduces structural bias, representing a critical area for advancing responsible AI. We address this gap by proposing SAFES – a Sequential PrivAcy and Fairness Enhancing data Synthesis procedure – which combines DP data synthesis with a fairness-aware pre-processing transformation. To our knowledge, this is the first work to attempt such an approach. The output of SAFES is a synthetic dataset that both has theoretical privacy guarantees and has been adjusted in a way to improve structural bias, which in turn improves fairness metrics for downstream classifiers.
SAFES has several benefits. First, it is fully tunable with regards to both the privacy guarantees and fairness constraints. Second, for tight fairness constraints, SAFES exhibits fairness robustness measured by various metrics across a wide range of privacy guarantees per our empirical results, implying that one can adjust the balance between privacy and utility without a significant sacrifice in fairness. Third, though our examples and experiments focus on a commonly used DP data synthesizer and a well-known fairness-aware data transformation methods, SAFES is a general framework and can admit different DP synthesizers of various DP guarantees and different fairness-aware data transformations satisfying various fairness metrics.
2 Related works
2.1 Data synthesis with formal privacy guarantees
Among the concepts developed for privacy protection, differential privacy (DP) (Dwork et al., 2006b, ; Dwork et al., 2006a, ) has emerged as the state-of-the-art framework for guaranteeing privacy when releasing information from sensitive data. To ensure DP, carefully calibrated random noise is added to outputs before release. DP mechanisms generally incur a loss of utility as a trade-off, governed by tunable privacy loss parameters. Many mechanisms have been developed for training ML models (Chaudhuri et al.,, 2011; Abadi et al.,, 2016) that satisfy DP. However, the outputs of these methods are limited to the specific chosen task, and the privacy loss accumulates with each task executed, which limits the number of analyses that can be performed under a fixed privacy budget. This fact motivates the use of DP for synthesizing datasets to be privacy-preserving counterparts to an original, sensitive dataset. Data users can analyze DP-synthesized data as if it were the original data without additional privacy loss due to the immunity to post-processing property of DP. Popular DP data synthesis methods include marginal-based synthesizers for discrete data (Zhang et al.,, 2017; McKenna et al.,, 2019; Eugenio and Liu,, 2021; McKenna et al.,, 2021, 2022). They compute a set of marginals with DP noise and train a model based on the sanitized marginals, from which data are synthesized. Statistical models and deep generative models are also used for DP data synthesis (Zhang et al.,, 2018; Bowen and Liu,, 2020; National Academies of Sciences, Engineering, and Medicine,, 2024).
2.2 Fairness-aware methods
A substantial amount of research and applications on data and algorithmic fairness in the context of AI/ML are for binary classification tasks due to its prevalence of the problem in practice and well-defined fairness metric, among other considerations. We focus on fairness in binary classification in the paper, unless stated otherwise.
Methods for ensuring fairness include pre-processing, in-processing, and post-processing procedures. Pre-processing methods (Hajian and Domingo-Ferrer,, 2013; Calmon et al.,, 2017) transform the dataset to be more fair prior to analysis. In-processing (Kamishima et al.,, 2011; Fish et al.,, 2016) and post-processing (Kamiran et al.,, 2012; Hardt et al.,, 2016) methods, on the other hand, modify the training process and the ultimate model decisions, respectively. While each type of method takes a different approach, the ultimate goal of all of the fairness methods described in this paper is to ensure decisions made by ML classifiers, where these concerns often arise, are done in a fair way. Pre-processing methods include methods that flip the protected attribute and/or outcome label until the proportion of each group receiving each outcome is similar (Hajian and Domingo-Ferrer,, 2013), and methods that use an optimization problem to learn a randomized pre-processing transformation balancing discrimination, distortion, and utility (Calmon et al.,, 2017). The ultimate goal in each case is to remove structural bias from the dataset itself, which in turn improves the fairness of classifiers trained with the pre-processed data. These methods are flexible because they are agnostic to downstream analysis and learning tasks and methods. In-processing methods include incorporating a regularizer when training an ML or AI model to penalize over-reliance on a protected attribute (Kamishima et al.,, 2011) and shifting the classifier decision boundary for unprivileged groups (Fish et al.,, 2016). Post-processing methods include defining a decision rule for a probabilistic classifier that assigns members of the unprivileged group the favorable outcome if the certainty of the classifier’s decision is below a certain threshold (Kamiran et al.,, 2012), and modifying a previously learned classifier to be non-discriminatory by solving a linear program (Hardt et al.,, 2016). Like DP, the methods for achieving fairness also generally incur a loss in utility.
2.3 Interplay between privacy, fairness, and utility
The impacts of privacy on fairness and vice versa have also been studied, including relationship between classification fairness and DP (Dwork et al.,, 2012); whether DP guarantees are inequitably applied across groups, exacerbating unfairness (Ekstrand et al.,, 2018); and whether fair ML techniques can put underprivileged groups at a greater privacy risk (Chang and Shokri,, 2021). It is reported that DP synthetic data often magnify unfairness (Ganev et al.,, 2022) and fairness transformations unevenly distribute privacy risk to underprivileged groups (Chang and Shokri,, 2021). There also exists a three-way trade-off between privacy, fairness, and utility. It is claimed that it is impossible for a single mechanism to achieve both DP and fairness with non-trivial classification accuracy (Cummings et al.,, 2019; Agarwal,, 2021); Classifiers trained on synthesized images via DP generative adversarial networks (GANs) exhibit reduced utility without fairness improvements (Cheng et al.,, 2021).
For specific learning tasks with DP and fairness constraints, methods exist for empirical risk minimization (Ding et al.,, 2020), logistic regression (Xu et al.,, 2019), and stochastic gradient descent (Tran et al.,, 2021), among others. Undersampling a training dataset prior to DP synthesis may also produce better fairness metrics on downstream classification tasks (Bullwinkel et al.,, 2022). Modifications have also been considered to achieve DP for in- and post-processing fairness methods (Jagielski et al.,, 2019). Data synthesis with DP and “justifiable fairness” (Salimi et al.,, 2019) guarantees can be achieved for a marginal-based DP synthesizer (McKenna et al.,, 2021) by ensuring all directed paths between the protected attribute(s) and the response variable in the graphical model representation pass through a non-protected attribute (Pujol et al.,, 2023). however, since the fairness modification is entangled with data synthesis in this approach, it does not apply in the increasingly common scenario where DP synthetic data were already released without fairness considerations. In addition, the approach only achieves justifiable fairness, which is a binary condition (yes or no) and only one of many possible fairness definitions and does not guarantee fairness by other metrics, especially those with a continuous fairness parameter for tuning the trade-off between utility and fairness constraints.
3 Definitions and notations
We introduce, in this section, the definitions and notations employed in our approach on differential privacy and fairness.
3.1 Differential privacy
Differential privacy (DP) provides a theoretical framework for privacy by bounding the influence of a single individual in a dataset on outputs from the dataset. Let $d(D,D^{\prime})=1$ denote two neighboring datasets $D,D^{\prime}$ differing by one individual.
Definition 1 (($\varepsilon,\delta)$-differential privacy (Dwork et al., 2006a, ; Dwork et al., 2006b, )).
A randomized mechanism $\mathcal{M}$ is ($\varepsilon,\delta)$-differentially private if for all $S\subset\textnormal{Range}(\mathcal{M})$ and for any neighboring datasets $D,D^{\prime}$,
$P(\mathcal{M}(D)\in S)\leq e^{\varepsilon}P(\mathcal{M}(D^{\prime})\in S)+\delta.$ (1)
$\varepsilon>0$ and $\delta\in[0,1)$ are privacy loss or budget parameters. If $\delta=0$, $(\varepsilon,\delta)$-DP reduces to $\varepsilon$-DP.
Essentially, DP ensures that a mechanism output cannot differ “too much” based on the inclusion or exclusion of a single individual. The individual differing between the neighboring datasets is arbitrary, meaning that DP guarantees hold simultaneously for all members of the dataset. Smaller $\varepsilon$ corresponds to more privacy. $\delta$ is often interpreted as the probability that $\varepsilon$-DP fails and is usually of $o(1/\mbox{poly}(n))$, where $n$ is the data sample size. There are various extensions (Bun and Steinke,, 2016; Mironov,, 2017; Dong et al.,, 2019) to the original DP definition in Definition 1. We present one of these extensions – zero-Concentrated DP that is used in our experiments.
Definition 2 (Zero-concentrated DP (zCDP) (Bun and Steinke,, 2016)).
Let $D_{\alpha}$ be the Rényi divergence of order $\alpha$. A randomized mechanism $\mathcal{M}$ satisfies $\rho$-zCDP if for any neighboring datasets $D,D^{\prime}$ and any $\alpha\in(1,\infty)$,
$D_{\alpha}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))\leq\alpha\rho.$ (2)
Theorem 1 (Conversion of $\rho$-zCDP to $(\varepsilon,\delta)$-DP (Canonne et al.,, 2020)).
Let $\mathcal{M}$ be a mechanism satisfying $\rho$-zCDP. For any given $\varepsilon\geq 0$, $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP with $\delta=\min_{\alpha>1}\frac{\exp\{(\alpha-1)(\alpha\rho-\varepsilon)\}}{\alpha-1}\left(1-\frac{1}{\alpha}\right)^{\alpha}$.
To release an output from a function on $D$ with DP guarantees, the randomized mechanism $\mathcal{M}$ is calibrated according to the global sensitivity of the function, defined as follows (though originally defined using the $\ell_{1}$-norm (Dwork et al., 2006b, ), we use a more general definition).
Definition 3 ($\ell_{p}$-global sensitivity (GS) (Liu,, 2019)).
Let $f$ be a (potentially vector-valued) function of a dataset $D$. The $\ell_{p}$-GS of $f$ is
$\Delta_{p,f}=\max_{d(D,D^{\prime})=1}\left\lVert f(D)-f(D^{\prime})\right\rVert_{p}.$ (3)
The GS of a function is the maximum difference in the function outputs on two neighboring datasets. The larger the GS is, the more noise is needed to achieved a fixed level of privacy guarantee via $\mathcal{M}$. The Gaussian mechanism (Definition 4) and the exponential mechanism (Definition 5) are two commonly used DP mechanisms.
Definition 4 (Gaussian mechanism (Dwork et al., 2006a, ; Bun and Steinke,, 2016)).
Let $D\in\mathcal{D}$ be a dataset. For a given function $f:\mathcal{D}\rightarrow\mathbb{R}^{n}$ with $\ell_{2}$-global sensitivity $\Delta_{2,f}$, the Gaussian mechanism is $\mathcal{M}(D)=f(D)+\mathbf{e}$, where $e_{i}$ is drawn independently from Gaussian distribution $\mathcal{N}(0,\sigma^{2})$.
Definition 5 (Exponential mechanism (McSherry and Talwar,, 2007)).
Let $\xi>0$ be a privacy loss parameter. For a given target function $f:\mathcal{D}\rightarrow\mathcal{R}$ and utility function $u:\mathcal{D}\times\mathcal{R}\rightarrow\mathbb{R}$ with $\ell_{1}$-global sensitivity $\Delta_{1,u}$, the exponential mechanism releases $r\in\mathcal{R}$ with probability $P(\mathcal{M}(D)=r)\propto\exp\left(\frac{\xi u(D,r)}{2\Delta_{1,u}}\right)$.
The Gaussian mechanism in Definition 4 achieves $(\varepsilon,\delta)$-DP for $\sigma^{2}=2\log(1.25/\delta)\Delta_{2,f}^{2}/\varepsilon^{2}$ and $\rho$-zCDP for $\sigma^{2}=\Delta_{2,f}^{2}/2\rho$. The exponential mechanism in Definition 5 achieves $\varepsilon$-DP for $\xi=\varepsilon$ and $\rho$-zCDP for $\xi=2\sqrt{2\rho}$ (Cesar and Rogers,, 2021).
Both $(\varepsilon,\delta)$-DP and $\rho$-zCDP are composable under repeated applications of randomized mechanisms to the same data. Let $\mathcal{D}$ be the space of all possible datasets $D$ and let $\mathcal{M}_{1}:\mathcal{D}\rightarrow\mathcal{R}_{1}$ and $\mathcal{M}_{2}:\mathcal{D}\times\mathcal{R}_{1}\rightarrow\mathcal{R}_{2}$ satisfy $(\varepsilon_{1},\delta_{1})$-DP ($\rho_{1}$-zCDP) and $(\varepsilon_{2},\delta_{2})$-DP ($\rho_{2}$-zCDP), respectively. Then $\mathcal{M}_{2}(D,\mathcal{M}_{1}(D))$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP ($(\rho_{1}+\rho_{2})$-zCDP). For $(\varepsilon,\delta)$-DP, this is known as basic composition (McSherry,, 2009). There exists an advanced composition theorem for DP that gives a tighter privacy loss bound (Dwork et al.,, 2010), but zCDP is still tighter (Bun and Steinke,, 2016). This motivates the practice of specifying the overall privacy loss in terms of $(\varepsilon,\delta)$-DP, applying Theorem 1 to convert it to $\rho$-zCDP, and using DP procedures of $\rho$-zCDP to compose privacy loss.
Another useful property of DP, in addition to privacy loss composition above, is its immunity to post-processing (Dwork et al., 2006b, ; Dwork et al., 2006a, ; Bun and Steinke,, 2016). That is, a post-processing procedure on an output of a DP mechanism, whether satisfying $(\varepsilon,\delta)$-DP or $\rho$-zCDP, does not incur further privacy loss, as long as the procedure does not access the original data. In the case of synthetic data generated from a DP synthesizer, any tasks done on the synthetic data would also satisfy the same DP guarantees as the synthetic data itself.
3.2 Fairness
We separate a dataset $D$ into three disjoint sets of variables, $D=(X,G,Y)$. $Y$ is a binary response variable consisting of a “favorable” outcome (e.g., approved for a loan) and an “unfavorable” outcome. The protected attributes $G$ distinguish different groups ($g_{1},g_{2},\ldots$) (e.g., race, gender). We assume each group can be collapsed to binary attributes representing a “privileged” and an “unprivileged” group. WLOG, we assume that the favorable outcome and the privileged group are each encoded as $1$. $X$ contains the non-protected predictors. Our ultimate fairness goal is to ensure that downstream classifiers trained on a dataset are fair with respect to the protected attributes. That said, it is also useful to measure the real-world bias present in the dataset itself. We introduce several metrics for measuring the dataset bias and downstream classifier fairness.
We consider a dataset to have structural bias if the probability of an observation receiving the favorable outcome is different for observations in the privileged and unprivileged groups. This can be measured by the conditional outcome difference (COD). By definition, COD is a property of the dataset rather than of an AI or ML algorithm. When $Y$ is independent of $G$, COD = 0, which represents the least biased dataset possible by this definition.
Definition 6 (Conditional outcome difference (COD)).
Given a dataset $D\!=\!(X,G,Y)$, let $g$ be a protected attribute in $G$. Then
$\textnormal{COD}(D)=P(Y=1|g=0)-P(Y=1|g=1).$ (4)
In statistical terms, structural bias is the same as the existence of dependency or correlation between group (e.g., privileged vs. unprivileged) and the outcome variable. Of course, not all such correlations indicate bias; some reflect genuine relationships. Here, we focus on cases where this correlation, if it exists, is not causal and would disappear if we conditioned on other relevant covariates that are not related to privilege (e.g. years of schooling). Additionally, for datasets without structural bias, imbalances in data representation between privileged and unprivileged groups can increase the likelihood of prediction bias in downstream ML algorithms. This prediction bias can be assessed using algorithmic fairness metrics, presented below.
Two commonly used fairness metrics for classifiers are statistical parity (Dwork et al.,, 2012), which ensures privileged and unprivileged groups are equally likely to get the favorable decision, and equalized odds (Hardt et al.,, 2016), which ensures privileged and unprivileged groups have identical true and false positive rates.
Definition 7 (Statistical parity difference (SPD) and average odds difference (AOD)(Dwork et al.,, 2012; Hardt et al.,, 2016)).
Given a dataset $D=(X,G,Y)$, let $g$ be a protected attribute from $G$. Let $\hat{Y}$ be the decision of a classifier learned from $D$. Then
$\textnormal{SPD}(D,\hat{Y})=P(\hat{Y}$ $=1|g=0)-P(\hat{Y}=1|g=1),$ (5)
$\textnormal{AOD}(D,\hat{Y})=0.5\big{[}$ $\big{(}P(\hat{Y}=1|Y=0,g=0)-P(\hat{Y}=1|Y=0,g=1)\big{)}+$
$\big{(}P(\hat{Y}=1|Y=1,g=0)-P(\hat{Y}=1|Y=1,g=1)\big{)}\big{]}.$ (6)
We define the conditional utility difference (CUD) to measure the balance in the utility of a classifier between the unprivileged and privileged groups. The utility function $u$ in CUD can be defined in various ways and is context-based. For example, $u(Y,\hat{Y}|g)$ can be the classification accuracy in group $g$.
Definition 8 (conditional utility difference (CUD)).
Given a dataset $D=(X,G,Y)$, let $g$ be a protected attribute from $G$ and $u(Y,\hat{Y}|g)$ be an arbitrary conditional utility function. Then
$\textnormal{CUD}(D,\hat{Y})=u(Y,\hat{Y}|g=0)-u(Y,\hat{Y}|g=1).$ (7)
CUD encompasses many fairness metrics in the literature like overall accuracy equality (Berk et al.,, 2021), predictive parity (Chouldechova,, 2017), and false positive/negative rate balance (Chouldechova,, 2017). For example, the CUD for false negative rate would use $u(Y,\hat{Y}|g)=P(\hat{Y}=0|Y=1,g)$.
For all fairness metrics presented in this section, a value closer to 00 is more fair. Negative values typically indicate unfairness in favor of the privileged group (e.g., a greater proportion of the privileged group is approved for a loan), while positive values favor the unprivileged group. We estimate each of these metrics empirically in our experiments.
4 The SAFES procedure
In this section, we first present SAFES – Sequential PrivAcy and Fairness Enhancing data Synthesis, and then briefly summarize the DP data synthesizer and the fairness-aware data transformation procedure used in our experiments, along with the reasons for selecting these methods for the experiments.
4.1 The algorithm
The SAFES algorithm consists of two sequential steps: DP data synthesis followed by fairness-aware data transformation. We provide the SAFES procedure in Algorithm 1, along with a formal claim for its DP guarantees.
Proposition 1.
SAFES satisfies DP avspacet privacy loss $\Theta_{1}$.
The proof is straightforward. Since $S$ satisfies DP with parameters $\Theta_{1}$, $D^{*}$ achieves $\Theta_{1}$-DP guarantees. Since the fairness data transformation $T$ only operates on $D^{*}$ and never accesses the original data $D$, the post-processing theorem for DP ensures the same DP guarantees on $\tilde{D}^{*}$.
Input : Dataset $D$, privacy loss parameters $\Theta_{1}$, fairness parameters $\Theta_{2}$
Output : Privacy-preserving and fairness-aware synthetic dataset $\tilde{D}^{*}$
1 Generate a privacy-preserving synthetic dataset $D^{*}$ at privacy loss $\Theta_{1}$ via a DP synthesizer $S$ (AIM in our experiments).
2 Transform $D^{*}$ with fairness parameters $\Theta_{2}$ to obtain $\tilde{D}^{*}$ via a fairness pre-processor $T$ (TOT in our experiments).
return $\tilde{D}^{*}$
Algorithm 1 The SAFES Procedure
Algorithm 1 accommodates any DP types. For example, if $\rho$-zCDP is used, then $\Theta_{1}=\{\rho\}$; if $(\varepsilon,\delta)$-DP or $\varepsilon$-DP is used, then $\Theta_{1}=\{\varepsilon,\delta\}$ or $\Theta_{1}=\{\varepsilon\}$. Algorithm 1 also permits any legitimate DP synthesizers described in the Introduction section. Similarly, any fairness-aware data pre-processor $T$ can be applied. The immunity to post-processing property of DP permits tuning $\eta$ to achieve the desired fairness-utility trade-off without additional privacy cost, since SAFES applies privacy and fairness sequentially. Figure 1 suggests the SAFES procedure is flexible and modular, allowing users to selectively apply its steps without requiring a strict start-to-finish approach. Users can directly release the DP synthetic data if fairness is not a concern or skip the DP synthesis step to work with previously released DP synthetic data.
Figure 1: The SAFES procedure and its applications
Algorithm 1 is general to permit any DP data synthesizer and fairness data pre-processor. Our experiments use AIM (McKenna et al.,, 2022) as the DP data synthesizer and a data transformation procedure that transforms $(X,Y)\subset D$ but leaves $G$ unchanged (Calmon et al.,, 2017) as the fairness-aware transformation step, respectively. For brevity, we refer to the latter simply as Triple-cOnstrained Transformation (TOT).
We chose AIM as the DP data synthesizer in the experiments because it has been shown to consistently outperform other synthesis methods in various utility metrics in previous works on DP data synthesis (McKenna et al.,, 2022). In addition, previous studies suggest that marginal-based methods, to which AIM belong, at large privacy budgets do not exacerbate group unfairness as much as deep generative model-based DP synthesis methods (Pereira et al.,, 2024). We chose TOT over other (limited) fairness-aware pre-processing methods (Hajian and Domingo-Ferrer,, 2013) because, as pointed out in (Calmon et al.,, 2017), TOT permits explicit control over the balance between group and individual fairness and also connects well with the broader statistical learning framework due to its probabilistic transformation.
4.2 Marginal-based DP data synthesis procedure
In general, for marginal-based synthesizers, a “workload” of $k$ marginals $W=\{\boldsymbol{\mu}^{\prime}_{1},\boldsymbol{\mu}^{\prime}_{2},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$ is first specified, with each $\boldsymbol{\mu}^{\prime}_{i}$ representing a marginal on data $D$ (e.g., a 2-way marginal on race and sex). A set of $m$ marginals from $W$ are then selected (possibly with replacement) and sanitized via DP mechanisms to compute the output marginals $\hat{\boldsymbol{\mu}}=\{\hat{\boldsymbol{\mu}}_{1},\ldots,\hat{\boldsymbol{\mu}}_{m}\}$. From the estimated DP marginal counts, a distribution $\hat{p}$ is learned that minimizes the difference between its marginals and $\hat{\boldsymbol{\mu}}$. $\hat{p}$ can then be sampled to generate synthetic data $D^{*}$ with DP guarantees.
AIM, the method used for the DP data synthesis subroutine of the SAFES procedure in our experiments, makes several improvements to this standard procedure, as listed in Algorithm 2.
Input : Dataset $D=(X,G,Y)$ with $d$ variables, marginal workload $W=\{\boldsymbol{\mu}^{\prime}_{1},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$, privacy loss $\rho$
Output : Privacy-preserving synthetic dataset $D^{*}$
Initialize $\sigma_{0}$, $\xi_{0}$ // conservative initialization is recommended
1 $\hat{\boldsymbol{\mu}}_{0}\leftarrow\boldsymbol{\mu}_{0}+\mathcal{N}(\mathbf{0},\sigma_{0}^{2}I)$, where $\boldsymbol{\mu}_{0}$ contains all 1-way marginals on $D$
2 $\hat{p}_{0}\leftarrow\arg\min_{p}\sum_{i=1}^{d}\left\lVert\boldsymbol{\mu}_{0,p}[i]-\hat{\boldsymbol{\mu}}_{0}[i]\right\rVert_{2}^{2}/\sigma_{0}$, where $\boldsymbol{\mu}_{0,p}$ contains all 1-way marginals based on distribution $p$ for data $D$
3 $\rho_{\textnormal{used}}\leftarrow d/(2\sigma^{2}_{0})$
for $j=1$ to $k$ do $w_{i}\leftarrow\sum_{i=1}^{k}|\boldsymbol{\mu}^{\prime}_{j}\cap\boldsymbol{\mu}^{\prime}_{i}|$ end // ||$\cdot$|| ⋅ | is the cardinality
4 $t\leftarrow 0$; $\sigma_{t+1}\leftarrow\sigma_{0}$; $\xi_{t+1}\leftarrow\xi_{0}$
5 while $\rho_{\textnormal{used}}<\rho$ do
6       $t\leftarrow t+1$
       $\rho_{\textnormal{used}}\leftarrow\rho_{\textnormal{used}}+\xi_{t}^{2}/8+1/(2\sigma^{2}_{t})$ // privacy loss accounting
       Determine a computationally feasible set of marginals $W^{\prime}\subseteq W$ and select $\boldsymbol{\mu}^{\prime}_{j}\in W^{\prime}$ via the exponential mechanism with $u(D,\boldsymbol{\mu}^{\prime}_{j})\!=\!w_{j}(\!\|\boldsymbol{\mu}^{\prime}_{j}\!-\!\boldsymbol{\mu}^{\prime}_{j,\hat{p}_{t-1}}\!\|_{1}\!\!-\!(2/\pi)^{1/2}\sigma_{t}n_{\boldsymbol{\mu}^{\prime}_{j}}\!)$ at privacy loss $\xi_{t}$, where $\boldsymbol{\mu}^{\prime}_{j,\hat{p}_{t-1}}$ is the selected marginal but measured by $\hat{p}_{t-1}$ // $u$ favors marginals with larger improvement in expected error under $\hat{p}_{t-1}$ and higher-order marginals; $n_{\boldsymbol{\mu}^{\prime}_{j}}$ is the number of cells in marginal $\boldsymbol{\mu}^{\prime}_{j}$
7       $\hat{\boldsymbol{\mu}}_{t}\leftarrow\boldsymbol{\mu}^{\prime}_{j}+\mathcal{N}(\mathbf{0},\sigma_{t}^{2}I)$
       $\hat{p}_{t}\leftarrow\arg\min_{p}\sum_{i=0}^{t}\left\lVert\boldsymbol{\mu}_{i,p}-\hat{\boldsymbol{\mu}}_{i}\right\rVert_{2}^{2}/\sigma_{i}$ // graphical model learning
      Update $\sigma_{t+1}$ and $\xi_{t+1}$ // according to Algorithm 3 of AIM arXiv paper (McKenna et al.,, 2022)
8
9 end while
10Sample $D^{*}=(X^{*},G^{*},Y^{*})$ from $\hat{p}_{t}$
11 return $D^{*}$
Algorithm 2 the AIM procedure for $\rho$-zCDP guarantees(McKenna et al.,, 2022)
For illustration, the steps in Algorithm 2 are listed with $\rho$-zCDP guarantees, but other DP types with the associated mechanisms can be used. The user first provides a workload $W$ to be considered for learning/updating a DP graphical model. The initial graphical model $\hat{p}_{0}$ is formed from all one-way marginals, then adaptively updated with higher-order marginals selected with replacement from $W$ in order to optimize the utility of the graphical model. The privacy budget spending scheme and update rule ($\sigma_{t+1}$ and $\xi_{t+1}$ on line 13 of Algorithm 2) ensure that marginals measured in later iterations receive a greater portion of the privacy budget when necessary to ensure useful marginal measurements. In particular, if the difference between the marginals produced by $\hat{p}_{i}$ and $\hat{p}_{i-1}$ is small, meaning little information is gained this round, $\xi_{i+1}=2\xi_{i}$ and $\sigma_{i+1}=\sigma_{i}/2$. See Algorithm 3 in the arXiv version of the AIM paper (McKenna et al.,, 2022) for more details. Considerations are also taken to ensure computational tractability for the graphical representation. These iterations proceed until the privacy budget is exhausted, at which point the current graphical model is sampled to obtain the synthetic dataset.
Since AIM only works with categorical or ordinal variables, for continuous variables, some discretization will be needed before applying AIM.
4.3 Fairness-aware data transformation
The TOT method transforms $(X,Y)$ to $(\tilde{X},\tilde{Y})$, but leaves $G$ unchanged. We summarize the main idea of TOT below and refer readers to the original paper (Calmon et al.,, 2017) for more details and theoretical results. TOT learns a randomized mapping $T$ in the form of a conditional distribution $q_{\tilde{X},\tilde{Y}|X,G,Y}$ by solving an optimization problem that balances controlling discrimination between groups (group fairness), limiting distortion of individual observations (individual fairness), and maintaining a data distribution similar to the one before the transformation (utility preservation). For discrimination control, it requires $q_{\tilde{Y}|G}$ be similar for any two groups. Specifically,
$J(q_{\tilde{Y}|G}(y|g_{1}),q_{\tilde{Y}|G}(y|g_{2}))=|q_{\tilde{Y}|G}(y|g_{1})/q_{\tilde{Y}|G}(y|g_{2})-1|\leq\eta_{y,g_{1},g_{2}}$ (8)
for all $y$ and groups $g_{1},g_{2}\in G$. $\eta$ controls the trade-off between enforcing group fairness and maintaining statistical relationships in the input dataset; different $\eta$ can be used for different groups or response values. We denote $\eta_{y,g_{1},g_{2}}$ for different $\{y,g_{1},g_{2}\}$ collectively by $\boldsymbol{\eta}$. Enforcing group fairness via equation (8) risks changing certain individuals unrealistically. To balance group and individual fairness, TOT limits the conditional expected distortion. Let $\phi:(\mathcal{X}\times\mathcal{Y})^{2}\rightarrow\mathbb{R}$ be a nonnegative distortion function, where $\mathcal{X}$ and $\mathcal{Y}$ are the domains of $X$ and $Y$. TOT requires
$\mathbb{E}\left[\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})|G=\mathbf{g},X=\mathbf{x},Y=y\right]\leq c_{\mathbf{g},\mathbf{x},y}$ (9)
for all $(\mathbf{g},\mathbf{x},y)$. If $\phi$ is binary-valued, equation (9) reduces to the probability of an undesirable mapping
$P\left(\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=1|G=\mathbf{g},X=\mathbf{x},Y=y\right)\leq c_{\mathbf{g},\mathbf{x},y}.$ (10)
Similar to $\eta$, different values of $c_{\mathbf{g},\mathbf{x},y}$ can be used for different groups or response values. We denote the bounds $c_{\mathbf{g},\mathbf{x},y}$ collectively as $\mathbf{c}$. The distortion function is also customizable. This customization should be based on the context of the problem, and there is not necessarily a single “best” function for a given case. For example, a mapping that modifies an individual’s age by several decades should receive a high distortion value $\phi$, while a mapping that does not change an individual’s age would receive a 00 distortion value. Finally, to ensure the transformation does not drastically modify the distribution of the input dataset, the objective function minimizes the total variation distance (TVD), which, for discrete attributes, is
$\textnormal{TVD}(q_{\tilde{X},\tilde{Y}},q_{X,Y})=\textstyle\frac{1}{2}\sum_{\mathbf{x},y}\left|q_{\tilde{X},\tilde{Y}}(\mathbf{x},y)-q_{X,Y}(\mathbf{x},y)\right|.$ (11)
Equations (8), (9), and (11), plus the constraint that $q_{\tilde{X},\tilde{Y}|D}$ is a valid distribution, give the optimization
$\underset{q_{\tilde{X},\tilde{Y}|D}}{\text{minimize}}\textstyle\frac{1}{2}\sum_{\mathbf{x},y}\left|q_{\tilde{X},\tilde{Y}}(\mathbf{x},y)-q_{X,Y}(\mathbf{x},y)\right|,\;\text{subject to}$ (12)
$J(q_{\tilde{Y}|G}(y|g_{1}),q_{\tilde{Y}|G}(y|g_{2}))\leq\eta_{y,g_{1},g_{2}},$ (13)
$\mathbb{E}\left[\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})|G=\mathbf{g},X=\mathbf{x},Y=y\right]\leq c_{\mathbf{g},\mathbf{x},y},$
which can be solved via a standard convex solver like the embedded conic solver (ECOS) (Domahidi et al.,, 2013) available in the CVXPY library (Diamond and Boyd,, 2016).
5 Experiments
We run two experiments using the SAFES procedure on the Adult (also known as the 1994 US “Census Income”) (Becker and Kohavi,, 1996) and the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) (ProPublica,, 2016) datasets. Each of these datasets are publicly available for download. We evaluate the three-way privacy-fairness-utility trade-offs in SAFES against data synthesis or transformation methods that focus on either privacy preservation or fairness enhance, but not both, as well as against a baseline with no privacy or fairness enhancing procedure applied.
5.1 Data
The Adult dataset is a subset of US 1994 Census income data on 48,842 individuals. The dataset is frequently used in both privacy and fairness literature due to the inclusion of sensitive variables (e.g., income) and the encoding of real-world discrimination (e.g., pay disparities based on gender/race). We consider a subset of 4 features (race, sex, education, age) and one response variable $Y$ (income). The favorable outcome for $Y$ is $Y\geq\$50$. The protected attributes are $G$ = {race, sex} with “white” and “male” being the privileged groups, respectively. These variables were preprocessed as follows. All non-white races were collapsed into one “non-white” category, age was compressed into decades, and education below 11th grade and above a Bachelor’s degree were combined into one category each. Table 1 summarizes the variables.
Feature Levels
Race {White, Non-white}
Sex {Male, Female}
Age (years) {17-26, 27-36, …, 87-96}
Education {< 11th grade, 11th grade, High school, Some college,
Associate’s, Vocational, Bachelor’s, Graduate}
[…]
Table 1: Variables in the experiment on the Adult dataset.
The COMPAS dataset contains criminal history information for 6,172 defendants. While it is more commonly used in fairness literature, it also carries privacy concerns; even inferring membership in the dataset is equivalent to disclosing that an individual was accused of a crime. We consider a subset of five feature variables (race, sex, age, number of priors, and degree of charge) and one response variable $Y$ (recidivism within two years) on a subset of 5,278 “African-American” and “Caucasian” individuals. The favorable outcome for $Y$ is that the individual did not recidivate. The protected attributes are $G$ = {race, sex} with “Caucasian” and “female” being the privileged groups, respectively. Table 2 summarizes the variables. Similarly, we discretized age and number of priors into 3 categories.
Feature Levels
Race {Caucasian, African-American}
Sex {Male, Female}
Age {<25, 25-45, >45}
Charge degree {Felony, Misdemeanor}
Number of priors {0, 1-3, >3}
Recidivism {Yes, No}
Table 2: Variables in the experiment on the COMPAS dataset.
5.2 Experiment and Algorithmic settings
For the AIM method, we set the workload of marginals $W$ to be all two-way marginals in both experiments. Though Algorithm 2 uses $\rho$-zCDP, our results are presented using $(\varepsilon,\delta)$-DP with $\delta=10^{-9}$ (which is $o(1/n)$) via the conversion theorem. This facilitates comparisons with other works, as $(\varepsilon,\delta)$-DP is the most common type of DP used in the literature.
In the Adult experiment, we set $\phi_{A}(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=$ 3 in equations (9) and (10) if education changes by more than one stage or if age changes by more than a decade, 2 if age changes by a decade, 1 if income decreases, and 0 otherwise. If more than one condition is satisfied, $\phi_{A}$ outputs the largest possible result.
In the COMPAS experiment, we define individual distortion functions $\phi_{i}$ for each variable (e.g., $\phi_{\textnormal{Age}}$) that all output 0 for no change. Additionally, $\phi_{\textnormal{Age}}$ and $\phi_{\textnormal{Priors}}$ output 1 if the value is changed to the adjacent category and 2 if changed to a non-adjacent category (e.g., < 25 to > 45 for age). $\phi_{\textnormal{Charge}}$ outputs 1 if the value is changed from felony to misdemeanor and 2 if changed from misdemeanor to felony. $\phi_{\textnormal{Recidivism}}$ outputs 2 if the value is changed. We define the distortion function $\phi_{C}$ for equations (9) and (10) as the sum of the individual distortion functions.
In both experiments, we convert $\phi_{A}$ and $\phi_{C}$ into a binary $\phi$ using threshold values $(0.99,1.99,2.99)$. We set $\mathbf{c}=(0.1,0.05,0)$ corresponding to each threshold for the distortion condition (equation (10)), so that a few minor changes to each record are permissible with small probability, but large changes or changes to too many attributes are strongly discouraged. For example, in the Adult experiment, the learned fairness transformation $q$ decreases income with probability $0.1$ and changes age by at most a decade with probability $0.05$, but is not allowed to change education by more than one stage or age by more than one decade. Any other change is freely permitted. We also fix each $\eta_{y,g_{1},g_{2}}=\eta$ for a given value $\eta$.
In each experiment, we split the original data into a training set and a test set in an 80/20 ratio and encode the favorable result as 1 and the unfavorable result as 0. Synthetic data of the same sample size as the training data were generated via SAFES in Algorithm 1 using AIM as the DP data synthesizer and TOT for the fairness transformation. As the privacy- and fairness- aware transformations in SAFES are probabilistic, we run 35 repeats to summarize the average performance and stability. These repeats are performed at a list of different privacy budget parameter $\varepsilon$ and fairness parameter $\eta$ described in Table 3 (exception for the original dataset). The values of $\eta_{1}$ and $\eta_{2}$ were chosen to have some separation in the metric values among different settings. Other values could just as easily have been chosen, and we would expect a continuous transition of metric values between chosen $\boldsymbol{\eta}$ (i.e., $\eta=0.05$ would produce a line in Figure 3 somewhere in between the $\eta=0.025$ and $\eta=0.1$ lines).
original DP $\varepsilon\in\{10^{-2},10^{-1.5},\ldots,10^{1}\};\delta=10^{-9}$
original none privacy-preserving only
fairness $\eta\in\{\eta_{1},\eta_{2}\}$ fairness-aware only SAFES
Table 3: Privacy and fairness algorithmic settings. $(\eta_{1},\eta_{2})=(0.025,0.1)$ for Adult, $(0.08,0.15)$ for COMPAS. The workload $W$ in the AIM synthesizer contained all two-way marginals in both experiments.
We use the SmartNoise SDK (OpenDP,, 2023) for the AIM subroutine in the experiments, and the AIF360 library (Bellamy et al.,, 2018) for the TOT subroutine. The randomness associated with DP occasionally makes the fairness transformation infeasible for small $\varepsilon$ (e.g., $\approx 30\%$ of the attempts failed for $\varepsilon\in\{10^{-2},10^{-1.5}\}$ and $\eta=0.08$ for COMPAS). Therefore, even though 35 repeats were attempted for each combination of $\varepsilon$ and $\eta$, a small number of the results presented in Figures 3 and 4 are summarized based on $<35$ repeats.
To evaluate the utility of the synthetic data, we 1) compare the data with the original training data by measuring the TVD between all one-, two-, and three-way marginals for the datasets; 2) perform a Kolmogorov-Smirnov (KS) test to evaluate the distributional similarity between the original training and the synthetic datasets; and 3) measure, for a logistic regression classifier trained on the synthetic data, the prediction accuracy, false positive (FP) and false negative (FN) rates, F1-score, and the ROC AUC on the test data.
To evaluate fairness, we measure the COD of the synthetic data (to understand the change in structural bias) and the SPD, the AOC, and the FP and FN rate balance for the logistic classifier fairness. Each of these metrics are evaluated with the protected attribute being race and sex individually, as well as jointly (e.g., white male as the privileged group).
5.3 Results
Examples of graphical models obtained via Algorithm 2 for the Adult and COMPAS experiments are shown in Figure 2. All the general utility, classification performance, and fairness metrics yield similar insights. For this reason, as well as to conserve space, we present results for a representative sample of utility and fairness metrics in Figures 3 and 4. Complete numeric results and additional figures are found in the supplementary information.
Figure 2: Examples of graphical model representations learned by the AIM algorithm for the Adult (left) and COMPAS (right) examples. The Adult example was obtained with $\varepsilon=0.1$, while the COMPAS example was obtained with $\varepsilon=1$.
Figure 3: Examples of the privacy (different lines) vs fairness (y-axis) vs utility (x-axis) trade-off in the Adult experiment. In each plot, each point on a line represents the mean and the error bar indicates ±$\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; the smaller $x$-axis values, the better the utility.
Figure 4: Examples of the privacy (different lines) vs fairness (y-axis) vs utility (x-axis) trade-off in the COMPAS experiment. In each plot, each point on a line represents the mean and the error bar indicates ±$\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; the smaller $x$-axis values, the better the utility.
The observations from Figures 3 and 4 on the privacy-fairness-utility trade-off can be summarized as follows. In addition, observations (1) to (3) below can also be seen in Supplementary Figures S1-S3 (Adult) and S8-S10 (COMPAS), while observations (4) and (5) can be seen in Supplementary Figures S4-S7 (Adult) and S11-S14 (COMPAS).
  1. (1)
    As expected, stringent privacy (small $\varepsilon$) and fairness (small $\eta$) parameters result in higher general utility loss (larger marginal TVDs and KS test statistics) and worse classifier prediction performance (accuracy, F1-score, AUC, etc.).
  2. (2)
    For small $\varepsilon$, DP noise is more influential on utility than the fairness parameter $\eta$; for larger $\varepsilon$, $\eta$ has a more noticeable impact, especially on the dataset utility. The fairness transformation generally has little impact on the classifier prediction performance, with or without DP guarantees.
  3. (3)
    Even for strong privacy guarantees and fairness constraints ($\varepsilon\in(0.05,1)$ and small $\eta$), SAFES synthetic data have similar utility to the original by several metrics, including the TVD of three-way marginals, which were not explicitly sanitized in the privacy step of SAFES.
  4. (4)
    In general, at very small $\varepsilon$, the SAFES synthetic data display more fairness than the original largely due to the large amount of noise injected for the DP guarantees but have poor utility. For $\varepsilon$ large enough to avoid this but still small enough to yield strong privacy guarantees, the synthetic data show significantly enhanced fairness for all three protected groups compared to the original that favors the privileged group, along with similar utility to the original.
  5. (5)
    For small $\eta$, fairness guaranteed by SAFES is robust to changes in privacy loss, allowing confident calibration of the privacy-utility balance while maintaining fairness. Additionally, the achieved fairness stabilizes as $\eta$ decreases.
Figures 3 and 4 suggest the FN rate balance in several cases is positive rather than negative. At first glance, this seems to indicate that the “unprivileged” sex group (female in Adult, male in COMPAS) is more likely to have the favorable outcome per this metric. However, for this metric, a positive value is rather an indication that more unprivileged than privileged individuals are incorrectly classified to the unfavorable outcome, which is still a disadvantageous result for the unprivileged group. Additionally, when measuring COD jointly on race+sex in the COMPAS experiment, the COD of the DP-only synthetic data (solid blue line) does not appear to converge to the COD of the original dataset (solid blue dot). This is an artifact of the AIM subroutine in SAFES when selecting from only one- and two-way marginals, as we did in our evaluations. Three-way relationships, if they exist, such as between race, sex, and recidivism, may not be fully captured (see, Figure 2 in Section “Details of study method and experiment procedures” for an example). We also note that the classifier utility (accuracy, F1-score, and AUC) at the baseline without privacy and fairness procedure are lower than the state-of-the-art for these datasets, though they outperform a trivial classifier for each dataset. This is due in part to the known scalability limitations (McKenna et al.,, 2022; Hu et al.,, 2024) for both AIM and TOT that limit considering too many attributes when synthesizing data, resulting in less predictors are available to fit classifiers than in other applications. That said, the performance of our classifier are similar to those obtained in other works on DP data synthesis using these datasets. For example, the $\approx 0.8$ accuracy in the Adult experiment (Supplementary Fig. S3) matches the accuracy obtained by Pujol et al., (2023); the $\approx 0.7$ AUC in the COMPAS experiment (Supplementary Fig. S10) outperforms the AUC by Pereira et al., (2024). Finally, TOT seems to increase the F1-score in the Adult experiment, which is counter-intuitive. We conjecture this was a spurious effect and false positive signal due to the number of metrics analyzed, especially considering that no other metric in either experiment yield similar observation. Nevertheless, it may be worth more in depth investigation in future work.
6 Discussion
We have presented the SAFES procedure, which synthesizes data that simultaneously achieve DP guarantees and satisfy fairness constraints. We have performed experiments on two real datasets to evaluate the privacy, fairness, and utility trade-off in SAFES. The results clearly demonstrate that SAFES can be applied to real-world scenarios to synthesize and release datasets that satisfy both strong DP guarantees and improved downstream classifier fairness metrics, without significantly degrading the general utility of the dataset or limiting to specific downstream learning tasks. To implement SAFES in practice, we recommend choosing the smallest possible $\varepsilon$ that gives an acceptable utility metric for the target problem, then choosing the smallest possible $\eta$ that still permits a solvable optimization problem for the fairness transformation.
While SAFES is a general framework for all data types, we focused on the AIM DP synthesizer and the TOT fairness pre-processor that are limited to categorical and discretized numeric variables in the experiments. For both AIM and TOT, the number of marginals and marginal cell counts grow exponentially with the numbers of variables and categories in each variable; in addition, the number of constraints in the TOT optimization also increases with the data dimensionality. In future work, we will examine the applications of SAFES in datasets with mixed categorical and numeric variables, using more general statistical or deep generative models to generate synthetic data and developing new fairness data transformations permitting numeric data, though such attempts so far have not produced promising results (Pereira et al.,, 2024). We will also develop methods to improve the scalability of SAFES in high-dimensional settings. Another interesting direction for future research is to extend the SAFES framework to satisfy fairness constraints in settings other than binary classification, such as multi-class classification, regression, ranking, and clustering. One challenges in these extensions is a lack of a set of accepted meaningful and interpretable fairness criteria.
The SAFES procedure has the potential for positive societal impacts in a wide range of fields (e.g., healthcare, hiring, criminal justice), where privacy and fairness considerations are necessary in the deployment of responsible AI. Organizations might consider integrating SAFES into their data-driven decision making pipeline, confident that they are gaining useful insights from synthetic data with guaranteed privacy and improved fairness. On a cautionary note, SAFES should not be used as a black box. It is important to understand the privacy and fairness requirement for a given problem before implementing SAFES; a lack of understanding of these implications can easily result in unsubstantiated claims about privacy and/or fairness, which would likely exacerbate the issue.
Data and code
The Adult (Becker and Kohavi,, 1996) and COMPAS (ProPublica,, 2016) datasets used in the experiments in this paper are publicly available on the UCI Machine Learning Repository and Kaggle, respectively. The code for the experiments in this paper can be found at https://github.com/sgiddens/SAFES.
Acknowledgments
This work is supported by the University of Notre Dame Schmitt Fellowship and Lucy Graduate Scholarship to S.G. A portion of the work was carried out while S.G. was an intern at SandboxAQ, and S.G. acknowledges helpful discussion from supervisors Nicolas Gama and Sandra Guasch.
Supplementary Materials
Additional figures for the Adult and COMPAS experiments
We present several additional figures for both the Adult and COMPAS experiments. Supplementary Fig. 5 through Supplementary Fig. 11 are for the Adult experiment, while Supplementary Fig. 12 through Supplementary Fig. 18 are for the COMPAS experiment.
We note that the error bands for different values of $\eta$ in the fairness-only simulations can sometimes be fairly large, especially for the larger values of $\eta$ in our experiments. Larger $\eta$ mean more flexibility in the permitted dataset distortions, and hence greater variability in the learned randomized fairness transformation. Smaller $\eta$, on the other hand, would only permit very specific changes and thus have less variability. This is most noticeable in the $\eta=0.1$ case of the right-hand plot of Supplementary Fig. 6. However, in that case, the entire error band of p-values falls in the “fail-to-reject” region of the plot, so they all essentially provide the same information.
Figure 5: Mean $\pm$ 1 SD (error bars and shaded regions) summed TVD in each marginal set for 1-way, 2-way, and 3-way marginals between the synthetic data vs the original data for the Adult experiment.
Figure 6: Mean $\pm$ 1 SD (error bars and shaded regions) test statistic and corresponding p-value for the KS test comparing original and synthetic datasets for the Adult experiment. Statistical significance threshold of $\alpha=0.05$ is marked in red in the plot on the right.
Figure 7: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SPaFAS synthetic data for the Adult experiment.
Figure 8: Mean $\pm$ 1 SD (error bars and shaded regions) COD, measured with race, sex, and race+sex as the protected attribute, for the Adult experiment.
Figure 9: Mean $\pm$ 1 SD (error bars and shaded regions) SPD, with race, sex, and race+sex as the protected attribute, for the Adult experiment.
Figure 10: Mean $\pm$ 1 SD (error bars and shaded regions) FN rate balance, with race, sex, and race+sex as the protected attribute, for the Adult experiment.
Figure 11: Mean $\pm$ 1 SD (error bars and shaded regions) AOD, with race, sex, and race+sex as the protected attribute, for the Adult experiment.
Figure 12: Mean $\pm$ 1 SD (error bars and shaded regions) summed TVD in each marginal set for 1-way, 2-way, and 3-way marginals between the synthetic data vs the original data for the COMPAS experiment.
Figure 13: Mean $\pm$ 1 SD (error bars and shaded regions) test statistic and corresponding p-value for the KS test comparing original and synthetic datasets for the COMPAS experiment. Statistical significance threshold of $\alpha=0.05$ is marked in red in the plot on the right.
Figure 14: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SPaFAS synthetic data for the COMPAS experiment.
Figure 15: Mean $\pm$ 1 SD (error bars and shaded regions) COD, measured with race, sex, and race+sex as the protected attribute, for the COMPAS experiment.
Figure 16: Mean $\pm$ 1 SD (error bars and shaded regions) SPD, with race, sex, and race+sex as the protected attribute, for the COMPAS experiment.
Figure 17: Mean $\pm$ 1 SD (error bars and shaded regions) FN rate balance, with race, sex, and race+sex as the protected attribute, for the COMPAS experiment.
Figure 18: Mean $\pm$ 1 SD (error bars and shaded regions) AOD, with race, sex, and race+sex as the protected attribute, for the COMPAS experiment.
Tabular results for the Adult and COMPAS experiments
We present the full set of tabular results for all privacy and utility metrics for both datasets. Supplementary Table 4 shows the results of the Adult experiment for various $\varepsilon$ privacy parameters with no fairness transformation. Supplementary Table 5 and Supplementary Table 6 show the results of the Adult experiment for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.1$ and $\eta=0.025$, respectively. Supplementary Table 7 shows the results of the COMPAS experiment for various $\varepsilon$ privacy parameters with no fairness transformation. Supplementary Table 8 and Supplementary Table 9 show the results of the COMPAS experiment for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.15$ and $\eta=0.08$, respectively.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.344 (0.110) 0.091 (0.027) 0.030 (0.010) 0.012 (0.005) 0.006 (0.001) 0.004 (0.001) 0.003 (0.001) 0.000 (0)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.058 (0.083) 0.588 (0.211) 0.906 (0.153) 0.980 (0.060) 1.000 (0)
Cumulative 1-way TVD 0.557 (0.094) 0.164 (0.035) 0.055 (0.011) 0.022 (0.005) 0.007 (0.001) 0.003 (0.000) 0.001 (0.000) 0.000 (0)
Cumulative 2-way TVD 2.141 (0.311) 0.775 (0.107) 0.369 (0.058) 0.133 (0.017) 0.052 (0.007) 0.025 (0.002) 0.010 (0.001) 0.000 (0)
Cumulative 3-way TVD 3.070 (0.428) 1.272 (0.141) 0.701 (0.110) 0.267 (0.027) 0.139 (0.013) 0.100 (0.003) 0.082 (0.001) 0.000 (0)
COD (race) -0.025 (0.117) -0.057 (0.076) -0.065 (0.045) -0.091 (0.028) -0.099 (0.006) -0.099 (0.002) -0.099 (0.001) -0.099 (0)
COD (sex) -0.107 (0.115) -0.187 (0.070) -0.194 (0.021) -0.192 (0.009) -0.193 (0.004) -0.193 (0.001) -0.193 (0.000) -0.193 (0)
COD (race + sex) -0.093 (0.105) -0.163 (0.065) -0.172 (0.025) -0.179 (0.014) -0.184 (0.004) -0.184 (0.001) -0.184 (0.000) -0.183 (0)
Accuracy 0.747 (0.025) 0.770 (0.010) 0.789 (0.007) 0.795 (0.001) 0.796 (0.001) 0.796 (0.000) 0.796 (0.000) 0.796 (0)
F1-score 0.124 (0.173) 0.304 (0.128) 0.446 (0.063) 0.463 (0.010) 0.464 (0.006) 0.464 (0.002) 0.464 (0.002) 0.463 (0)
TP rate 0.100 (0.151) 0.222 (0.117) 0.352 (0.073) 0.359 (0.015) 0.358 (0.009) 0.359 (0.003) 0.358 (0.002) 0.356 (0)
TN rate 0.959 (0.058) 0.949 (0.033) 0.932 (0.023) 0.938 (0.005) 0.939 (0.003) 0.939 (0.001) 0.940 (0.001) 0.940 (0)
FN rate 0.900 (0.151) 0.778 (0.117) 0.648 (0.073) 0.641 (0.015) 0.642 (0.009) 0.641 (0.003) 0.642 (0.002) 0.644 (0)
FP rate 0.041 (0.058) 0.051 (0.033) 0.068 (0.023) 0.062 (0.005) 0.061 (0.003) 0.061 (0.001) 0.060 (0.001) 0.060 (0)
AUC 0.607 (0.104) 0.707 (0.076) 0.791 (0.027) 0.819 (0.002) 0.823 (0.001) 0.824 (0.000) 0.824 (0.000) 0.824 (0)
SPD (race) -0.010 (0.036) -0.041 (0.059) -0.071 (0.039) -0.081 (0.022) -0.077 (0.009) -0.073 (0.003) -0.072 (0.001) -0.072 (0)
SPD (sex) -0.046 (0.085) -0.119 (0.084) -0.178 (0.051) -0.189 (0.016) -0.183 (0.008) -0.182 (0.002) -0.182 (0.002) -0.180 (0)
SPD (race + sex) -0.038 (0.072) -0.108 (0.082) -0.165 (0.052) -0.179 (0.016) -0.171 (0.007) -0.168 (0.003) -0.168 (0.002) -0.166 (0)
AOD (race) 0.003 (0.045) -0.025 (0.093) -0.041 (0.061) -0.058 (0.043) -0.044 (0.019) -0.036 (0.005) -0.035 (0.002) -0.036 (0)
AOD (sex) -0.046 (0.099) -0.141 (0.108) -0.211 (0.067) -0.239 (0.023) -0.231 (0.012) -0.229 (0.003) -0.228 (0.002) -0.226 (0)
AOD (race + sex) -0.030 (0.073) -0.107 (0.097) -0.162 (0.064) -0.188 (0.024) -0.175 (0.010) -0.169 (0.003) -0.169 (0.002) -0.168 (0)
FN rate balance (race) -0.013 (0.064) 0.028 (0.153) 0.042 (0.099) 0.070 (0.075) 0.043 (0.034) 0.028 (0.008) 0.027 (0.003) 0.029 (0)
FN rate balance (sex) 0.060 (0.142) 0.215 (0.158) 0.330 (0.098) 0.386 (0.036) 0.375 (0.018) 0.372 (0.004) 0.371 (0.003) 0.368 (0)
FN rate balance (race + sex) 0.034 (0.096) 0.153 (0.140) 0.233 (0.090) 0.284 (0.037) 0.262 (0.016) 0.252 (0.005) 0.251 (0.003) 0.250 (0)
FP rate balance (race) -0.007 (0.030) -0.021 (0.037) -0.040 (0.027) -0.045 (0.012) -0.045 (0.005) -0.043 (0.002) -0.043 (0.001) -0.042 (0)
FP rate balance (sex) -0.031 (0.061) -0.067 (0.060) -0.092 (0.038) -0.092 (0.012) -0.087 (0.006) -0.086 (0.002) -0.086 (0.001) -0.084 (0)
FP rate balance (race + sex) -0.027 (0.054) -0.062 (0.059) -0.090 (0.041) -0.093 (0.013) -0.088 (0.006) -0.086 (0.002) -0.086 (0.002) -0.085 (0)
Table 4: Full set of simulation results for the Adult dataset for various values of the privacy parameter $\varepsilon$ without the fairness transformation. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 runs were performed in each case. The final column represents a run that is also without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.327 (0.119) 0.091 (0.031) 0.034 (0.009) 0.015 (0.005) 0.010 (0.004) 0.007 (0.002) 0.006 (0.003) 0.004 (0.003)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.026 (0.051) 0.183 (0.286) 0.423 (0.304) 0.518 (0.372) 0.795 (0.321)
Cumulative 1-way TVD 0.555 (0.119) 0.170 (0.040) 0.054 (0.013) 0.022 (0.006) 0.008 (0.002) 0.004 (0.001) 0.003 (0.001) 0.002 (0.001)
Cumulative 2-way TVD 2.119 (0.411) 0.813 (0.106) 0.405 (0.052) 0.177 (0.016) 0.109 (0.007) 0.086 (0.003) 0.075 (0.003) 0.070 (0.002)
Cumulative 3-way TVD 3.031 (0.555) 1.320 (0.141) 0.773 (0.106) 0.367 (0.023) 0.271 (0.010) 0.240 (0.006) 0.225 (0.005) 0.180 (0.004)
COD (race) -0.015 (0.038) -0.025 (0.029) -0.029 (0.022) -0.047 (0.014) -0.046 (0.004) -0.048 (0.004) -0.047 (0.003) -0.046 (0.003)
COD (sex) -0.033 (0.053) -0.087 (0.028) -0.096 (0.006) -0.093 (0.003) -0.093 (0.003) -0.093 (0.003) -0.093 (0.003) -0.093 (0.002)
COD (race + sex) -0.030 (0.044) -0.077 (0.022) -0.086 (0.008) -0.093 (0.005) -0.092 (0.002) -0.093 (0.002) -0.093 (0.003) -0.092 (0.002)
Accuracy 0.746 (0.017) 0.768 (0.013) 0.785 (0.009) 0.794 (0.001) 0.794 (0.001) 0.794 (0.001) 0.794 (0.000) 0.794 (0.000)
F1-score 0.131 (0.171) 0.315 (0.099) 0.407 (0.077) 0.468 (0.010) 0.471 (0.003) 0.471 (0.002) 0.471 (0.001) 0.471 (0.001)
TP rate 0.109 (0.156) 0.226 (0.090) 0.308 (0.083) 0.368 (0.014) 0.373 (0.005) 0.372 (0.003) 0.372 (0.002) 0.372 (0.002)
TN rate 0.954 (0.060) 0.946 (0.031) 0.940 (0.019) 0.933 (0.005) 0.932 (0.002) 0.932 (0.001) 0.932 (0.001) 0.932 (0.001)
FN rate 0.891 (0.156) 0.774 (0.090) 0.692 (0.083) 0.632 (0.014) 0.627 (0.005) 0.628 (0.003) 0.628 (0.002) 0.628 (0.002)
FP rate 0.046 (0.060) 0.054 (0.031) 0.060 (0.019) 0.067 (0.005) 0.068 (0.002) 0.068 (0.001) 0.068 (0.001) 0.068 (0.001)
AUC 0.613 (0.088) 0.717 (0.026) 0.777 (0.030) 0.811 (0.003) 0.813 (0.002) 0.813 (0.001) 0.814 (0.001) 0.813 (0.001)
SPD (race) -0.018 (0.040) -0.030 (0.027) -0.044 (0.022) -0.068 (0.017) -0.057 (0.010) -0.060 (0.009) -0.058 (0.009) -0.055 (0.008)
SPD (sex) -0.017 (0.053) -0.085 (0.072) -0.099 (0.041) -0.129 (0.009) -0.127 (0.004) -0.128 (0.003) -0.128 (0.002) -0.126 (0.002)
SPD (race + sex) -0.019 (0.047) -0.075 (0.063) -0.090 (0.038) -0.123 (0.010) -0.117 (0.005) -0.119 (0.004) -0.117 (0.003) -0.116 (0.003)
AOD (race) -0.009 (0.050) -0.005 (0.039) -0.005 (0.030) -0.031 (0.033) -0.011 (0.016) -0.015 (0.014) -0.011 (0.015) -0.009 (0.013)
AOD (sex) -0.008 (0.058) -0.084 (0.107) -0.087 (0.058) -0.122 (0.014) -0.116 (0.005) -0.118 (0.005) -0.118 (0.003) -0.116 (0.003)
AOD (race + sex) -0.009 (0.046) -0.058 (0.081) -0.058 (0.044) -0.093 (0.017) -0.080 (0.007) -0.083 (0.007) -0.081 (0.006) -0.080 (0.005)
FN rate balance (race) 0.007 (0.071) -0.004 (0.063) -0.013 (0.050) 0.025 (0.057) -0.009 (0.024) -0.003 (0.021) -0.010 (0.023) -0.012 (0.021)
FN rate balance (sex) 0.013 (0.079) 0.129 (0.166) 0.136 (0.093) 0.193 (0.022) 0.183 (0.007) 0.187 (0.009) 0.186 (0.004) 0.184 (0.004)
FN rate balance (race + sex) 0.011 (0.060) 0.079 (0.119) 0.079 (0.066) 0.131 (0.028) 0.109 (0.011) 0.114 (0.011) 0.111 (0.009) 0.110 (0.008)
FP rate balance (race) -0.012 (0.031) -0.015 (0.018) -0.022 (0.014) -0.037 (0.010) -0.031 (0.007) -0.033 (0.007) -0.032 (0.007) -0.029 (0.006)
FP rate balance (sex) -0.004 (0.040) -0.040 (0.049) -0.037 (0.024) -0.052 (0.007) -0.049 (0.003) -0.050 (0.003) -0.050 (0.002) -0.049 (0.002)
FP rate balance (race + sex) -0.006 (0.035) -0.036 (0.044) -0.037 (0.023) -0.055 (0.007) -0.051 (0.003) -0.052 (0.003) -0.051 (0.003) -0.050 (0.002)
Table 5: Full set of simulation results for the Adult dataset for $\eta=0.1$ and for various values of the privacy parameter $\varepsilon$. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 runs were performed in each case. The final column represents a run without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.350 (0.128) 0.086 (0.025) 0.047 (0.013) 0.029 (0.004) 0.029 (0.003) 0.028 (0.002) 0.028 (0.002) 0.027 (0.001)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000)
Cumulative 1-way TVD 0.593 (0.120) 0.183 (0.036) 0.086 (0.014) 0.056 (0.005) 0.046 (0.003) 0.045 (0.002) 0.044 (0.002) 0.042 (0.001)
Cumulative 2-way TVD 2.252 (0.429) 0.825 (0.101) 0.478 (0.056) 0.263 (0.013) 0.222 (0.009) 0.213 (0.006) 0.209 (0.006) 0.202 (0.004)
Cumulative 3-way TVD 3.204 (0.583) 1.348 (0.128) 0.871 (0.099) 0.483 (0.016) 0.418 (0.012) 0.398 (0.007) 0.392 (0.006) 0.365 (0.006)
COD (race) -0.003 (0.013) -0.004 (0.008) -0.011 (0.009) -0.018 (0.004) -0.017 (0.005) -0.018 (0.004) -0.018 (0.003) -0.018 (0.003)
COD (sex) -0.015 (0.013) -0.024 (0.003) -0.023 (0.003) -0.022 (0.003) -0.022 (0.002) -0.022 (0.003) -0.022 (0.003) -0.022 (0.003)
COD (race + sex) -0.013 (0.012) -0.021 (0.004) -0.023 (0.004) -0.025 (0.003) -0.025 (0.002) -0.025 (0.003) -0.025 (0.003) -0.025 (0.003)
Accuracy 0.731 (0.032) 0.761 (0.014) 0.776 (0.006) 0.783 (0.003) 0.785 (0.002) 0.786 (0.001) 0.786 (0.001) 0.786 (0.000)
F1-score 0.162 (0.158) 0.329 (0.046) 0.408 (0.070) 0.474 (0.018) 0.482 (0.003) 0.484 (0.002) 0.483 (0.002) 0.483 (0.001)
TP rate 0.134 (0.141) 0.241 (0.054) 0.322 (0.084) 0.397 (0.024) 0.407 (0.005) 0.407 (0.002) 0.405 (0.003) 0.405 (0.001)
TN rate 0.926 (0.073) 0.931 (0.028) 0.925 (0.024) 0.910 (0.007) 0.908 (0.003) 0.910 (0.001) 0.911 (0.001) 0.911 (0.001)
FN rate 0.866 (0.141) 0.759 (0.054) 0.678 (0.084) 0.603 (0.024) 0.593 (0.005) 0.593 (0.002) 0.595 (0.003) 0.595 (0.001)
FP rate 0.074 (0.073) 0.069 (0.028) 0.075 (0.024) 0.090 (0.007) 0.092 (0.003) 0.090 (0.001) 0.089 (0.001) 0.089 (0.001)
AUC 0.598 (0.087) 0.694 (0.028) 0.758 (0.029) 0.789 (0.004) 0.788 (0.002) 0.790 (0.002) 0.790 (0.002) 0.789 (0.001)
SPD (race) -0.018 (0.029) -0.031 (0.021) -0.037 (0.022) -0.064 (0.017) -0.047 (0.010) -0.043 (0.006) -0.042 (0.005) -0.040 (0.001)
SPD (sex) -0.007 (0.032) -0.029 (0.032) -0.040 (0.023) -0.060 (0.011) -0.061 (0.004) -0.064 (0.002) -0.063 (0.002) -0.063 (0.001)
SPD (race + sex) -0.010 (0.030) -0.030 (0.031) -0.042 (0.024) -0.066 (0.012) -0.063 (0.004) -0.063 (0.003) -0.062 (0.002) -0.061 (0.001)
AOD (race) -0.009 (0.029) -0.004 (0.025) 0.007 (0.033) -0.024 (0.033) 0.010 (0.023) 0.019 (0.014) 0.020 (0.012) 0.024 (0.001)
AOD (sex) 0.010 (0.046) 0.009 (0.035) 0.015 (0.029) 0.003 (0.014) 0.002 (0.004) -0.000 (0.002) 0.000 (0.002) 0.000 (0.002)
AOD (race + sex) 0.004 (0.037) 0.009 (0.031) 0.014 (0.027) -0.006 (0.017) 0.004 (0.009) 0.007 (0.006) 0.008 (0.005) 0.009 (0.001)
FN rate balance (race) 0.005 (0.039) -0.009 (0.038) -0.032 (0.057) 0.016 (0.057) -0.043 (0.042) -0.060 (0.024) -0.062 (0.022) -0.068 (0.001)
FN rate balance (sex) -0.016 (0.064) -0.014 (0.047) -0.024 (0.042) -0.004 (0.023) -0.003 (0.006) -0.000 (0.003) -0.001 (0.003) -0.000 (0.002)
FN rate balance (race + sex) -0.009 (0.050) -0.018 (0.039) -0.029 (0.040) 0.002 (0.028) -0.016 (0.016) -0.022 (0.010) -0.024 (0.009) -0.026 (0.002)
FP rate balance (race) -0.012 (0.023) -0.016 (0.017) -0.018 (0.014) -0.033 (0.010) -0.023 (0.005) -0.022 (0.003) -0.021 (0.002) -0.020 (0.001)
FP rate balance (sex) 0.003 (0.029) 0.004 (0.024) 0.007 (0.016) 0.001 (0.006) 0.001 (0.003) -0.001 (0.002) -0.000 (0.001) -0.000 (0.002)
FP rate balance (race + sex) -0.001 (0.026) -0.001 (0.025) -0.000 (0.015) -0.010 (0.006) -0.007 (0.002) -0.008 (0.002) -0.008 (0.001) -0.007 (0.001)
Table 6: Full set of simulation results for the Adult dataset for $\eta=0.025$ and for various values of the privacy parameter $\varepsilon$. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 runs were performed in each case. The final column represents a run without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.553 (0.152) 0.261 (0.077) 0.104 (0.031) 0.054 (0.016) 0.018 (0.006) 0.011 (0.003) 0.009 (0.002) 0.000 (0.000)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.010 (0.027) 0.580 (0.306) 0.924 (0.134) 0.979 (0.041) 1.000 (0.000)
Cumulative 1-way TVD 1.362 (0.385) 0.520 (0.151) 0.193 (0.053) 0.064 (0.019) 0.018 (0.005) 0.007 (0.002) 0.003 (0.001) 0.000 (0.000)
Cumulative 2-way TVD 5.262 (1.167) 2.308 (0.572) 1.039 (0.171) 0.515 (0.090) 0.175 (0.034) 0.066 (0.012) 0.023 (0.003) 0.000 (0.000)
Cumulative 3-way TVD 8.800 (1.609) 4.349 (0.968) 2.227 (0.292) 1.236 (0.218) 0.476 (0.077) 0.266 (0.022) 0.210 (0.004) 0.000 (0.000)
COD (race) -0.017 (0.149) -0.061 (0.145) -0.029 (0.071) -0.085 (0.053) -0.114 (0.022) -0.126 (0.010) -0.130 (0.002) -0.129 (0.000)
COD (sex) 0.002 (0.279) 0.043 (0.279) -0.055 (0.095) -0.053 (0.076) -0.127 (0.042) -0.139 (0.014) -0.141 (0.006) -0.141 (0.000)
COD (race + sex) -0.018 (0.222) 0.013 (0.277) -0.067 (0.093) -0.100 (0.072) -0.178 (0.043) -0.189 (0.020) -0.191 (0.006) -0.136 (0.000)
Accuracy 0.512 (0.055) 0.517 (0.058) 0.581 (0.067) 0.649 (0.019) 0.670 (0.007) 0.675 (0.002) 0.675 (0.000) 0.675 (0.000)
F1-score 0.505 (0.261) 0.532 (0.208) 0.581 (0.169) 0.681 (0.033) 0.699 (0.014) 0.709 (0.004) 0.707 (0.001) 0.708 (0.000)
TP rate 0.618 (0.383) 0.602 (0.302) 0.589 (0.239) 0.682 (0.079) 0.694 (0.033) 0.714 (0.010) 0.706 (0.002) 0.709 (0.000)
TN rate 0.380 (0.396) 0.413 (0.313) 0.571 (0.222) 0.608 (0.087) 0.641 (0.028) 0.627 (0.010) 0.636 (0.003) 0.633 (0.000)
FN rate 0.382 (0.383) 0.398 (0.302) 0.411 (0.239) 0.318 (0.079) 0.306 (0.033) 0.286 (0.010) 0.294 (0.002) 0.291 (0.000)
FP rate 0.620 (0.396) 0.587 (0.313) 0.429 (0.222) 0.392 (0.087) 0.359 (0.028) 0.373 (0.010) 0.364 (0.003) 0.367 (0.000)
AUC 0.488 (0.071) 0.517 (0.067) 0.601 (0.083) 0.692 (0.023) 0.716 (0.003) 0.719 (0.001) 0.720 (0.001) 0.720 (0.000)
SPD (race) 0.015 (0.207) -0.066 (0.197) -0.218 (0.265) -0.268 (0.134) -0.270 (0.059) -0.270 (0.027) -0.284 (0.003) -0.280 (0.000)
SPD (sex) 0.018 (0.191) 0.063 (0.336) -0.191 (0.264) -0.210 (0.136) -0.264 (0.057) -0.247 (0.026) -0.251 (0.012) -0.268 (0.000)
SPD (race + sex) 0.023 (0.196) 0.037 (0.324) -0.234 (0.265) -0.271 (0.118) -0.288 (0.048) -0.278 (0.024) -0.286 (0.007) -0.287 (0.000)
AOD (race) 0.016 (0.203) -0.065 (0.199) -0.198 (0.268) -0.231 (0.137) -0.229 (0.059) -0.228 (0.028) -0.241 (0.004) -0.238 (0.000)
AOD (sex) 0.018 (0.197) 0.063 (0.335) -0.176 (0.272) -0.180 (0.142) -0.230 (0.059) -0.214 (0.027) -0.218 (0.013) -0.236 (0.000)
AOD (race + sex) 0.025 (0.197) 0.036 (0.326) -0.215 (0.274) -0.233 (0.127) -0.250 (0.050) -0.241 (0.027) -0.250 (0.007) -0.251 (0.000)
FN rate balance (race) -0.013 (0.205) 0.069 (0.193) 0.200 (0.259) 0.231 (0.135) 0.215 (0.067) 0.215 (0.029) 0.232 (0.003) 0.227 (0.000)
FN rate balance (sex) -0.017 (0.182) -0.073 (0.341) 0.166 (0.244) 0.175 (0.125) 0.217 (0.056) 0.195 (0.023) 0.200 (0.009) 0.213 (0.000)
FN rate balance (race + sex) -0.017 (0.186) -0.045 (0.318) 0.223 (0.241) 0.257 (0.103) 0.258 (0.046) 0.241 (0.018) 0.249 (0.005) 0.249 (0.000)
FP rate balance (race) 0.019 (0.204) -0.060 (0.215) -0.196 (0.281) -0.231 (0.142) -0.242 (0.051) -0.241 (0.027) -0.251 (0.004) -0.248 (0.000)
FP rate balance (sex) 0.018 (0.218) 0.053 (0.336) -0.187 (0.308) -0.184 (0.163) -0.244 (0.065) -0.233 (0.032) -0.235 (0.017) -0.259 (0.000)
FP rate balance (race + sex) 0.032 (0.221) 0.027 (0.346) -0.208 (0.318) -0.208 (0.160) -0.243 (0.059) -0.241 (0.036) -0.250 (0.009) -0.252 (0.000)
Table 7: Full set of simulation results for the COMPAS dataset for various values of the privacy parameter $\varepsilon$ without the fairness transformation. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 successful runs were performed in each case. The final column represents a run that is also without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.531 (0.136) 0.244 (0.081) 0.103 (0.029) 0.050 (0.012) 0.022 (0.006) 0.013 (0.003) 0.012 (0.003) 0.007 (0.002)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.009 (0.029) 0.380 (0.273) 0.845 (0.196) 0.870 (0.146) 0.997 (0.009)
Cumulative 1-way TVD 1.345 (0.296) 0.511 (0.176) 0.167 (0.042) 0.066 (0.021) 0.023 (0.006) 0.012 (0.003) 0.011 (0.003) 0.008 (0.002)
Cumulative 2-way TVD 5.180 (1.036) 2.292 (0.705) 1.010 (0.151) 0.528 (0.060) 0.210 (0.034) 0.111 (0.014) 0.084 (0.009) 0.071 (0.006)
Cumulative 3-way TVD 8.643 (1.535) 4.341 (1.157) 2.215 (0.267) 1.265 (0.145) 0.539 (0.077) 0.326 (0.033) 0.289 (0.019) 0.182 (0.013)
COD (race) -0.004 (0.063) -0.001 (0.043) -0.031 (0.055) -0.080 (0.046) -0.090 (0.020) -0.100 (0.006) -0.098 (0.005) -0.112 (0.004)
COD (sex) 0.002 (0.019) -0.004 (0.052) -0.012 (0.045) -0.045 (0.056) -0.087 (0.034) -0.092 (0.008) -0.095 (0.008) -0.095 (0.007)
COD (race + sex) 0.013 (0.045) -0.003 (0.060) -0.027 (0.054) -0.074 (0.044) -0.098 (0.023) -0.091 (0.015) -0.092 (0.014) -0.081 (0.010)
Accuracy 0.475 (0.060) 0.537 (0.061) 0.584 (0.061) 0.648 (0.018) 0.670 (0.007) 0.677 (0.004) 0.677 (0.004) 0.678 (0.001)
F1-score 0.333 (0.299) 0.550 (0.224) 0.630 (0.086) 0.684 (0.038) 0.702 (0.013) 0.712 (0.007) 0.712 (0.006) 0.714 (0.003)
TP rate 0.385 (0.396) 0.640 (0.333) 0.663 (0.170) 0.693 (0.085) 0.702 (0.031) 0.720 (0.016) 0.722 (0.013) 0.727 (0.007)
TN rate 0.587 (0.403) 0.409 (0.339) 0.487 (0.196) 0.593 (0.097) 0.630 (0.031) 0.623 (0.013) 0.620 (0.009) 0.617 (0.008)
FN rate 0.615 (0.396) 0.360 (0.333) 0.337 (0.170) 0.307 (0.085) 0.298 (0.031) 0.280 (0.016) 0.278 (0.013) 0.273 (0.007)
FP rate 0.413 (0.403) 0.591 (0.339) 0.513 (0.196) 0.407 (0.097) 0.370 (0.031) 0.377 (0.013) 0.380 (0.009) 0.383 (0.008)
AUC 0.500 (0.105) 0.552 (0.082) 0.604 (0.084) 0.694 (0.019) 0.714 (0.003) 0.718 (0.002) 0.718 (0.001) 0.719 (0.001)
SPD (race) 0.048 (0.127) -0.067 (0.175) -0.192 (0.316) -0.276 (0.114) -0.218 (0.060) -0.243 (0.018) -0.236 (0.016) -0.245 (0.012)
SPD (sex) 0.035 (0.155) -0.029 (0.130) -0.096 (0.200) -0.173 (0.075) -0.221 (0.057) -0.211 (0.019) -0.208 (0.015) -0.205 (0.010)
SPD (race + sex) 0.060 (0.149) -0.064 (0.123) -0.178 (0.269) -0.241 (0.076) -0.250 (0.037) -0.243 (0.019) -0.243 (0.015) -0.242 (0.010)
AOD (race) 0.045 (0.124) -0.062 (0.174) -0.172 (0.315) -0.238 (0.115) -0.176 (0.060) -0.199 (0.018) -0.193 (0.015) -0.202 (0.013)
AOD (sex) 0.032 (0.152) -0.023 (0.130) -0.079 (0.196) -0.141 (0.077) -0.185 (0.057) -0.174 (0.020) -0.171 (0.016) -0.169 (0.010)
AOD (race + sex) 0.056 (0.147) -0.058 (0.122) -0.153 (0.270) -0.198 (0.078) -0.208 (0.038) -0.200 (0.021) -0.200 (0.016) -0.200 (0.011)
FN rate balance (race) -0.046 (0.132) 0.057 (0.161) 0.183 (0.311) 0.246 (0.118) 0.164 (0.063) 0.190 (0.020) 0.183 (0.018) 0.192 (0.013)
FN rate balance (sex) -0.032 (0.167) 0.025 (0.119) 0.081 (0.191) 0.143 (0.072) 0.181 (0.059) 0.170 (0.021) 0.166 (0.017) 0.161 (0.009)
FN rate balance (race + sex) -0.061 (0.154) 0.061 (0.114) 0.184 (0.248) 0.241 (0.072) 0.230 (0.039) 0.223 (0.018) 0.221 (0.015) 0.216 (0.008)
FP rate balance (race) 0.043 (0.119) -0.067 (0.195) -0.161 (0.322) -0.230 (0.115) -0.188 (0.061) -0.208 (0.017) -0.203 (0.013) -0.211 (0.013)
FP rate balance (sex) 0.032 (0.141) -0.022 (0.147) -0.077 (0.209) -0.138 (0.085) -0.188 (0.057) -0.179 (0.022) -0.176 (0.017) -0.177 (0.012)
FP rate balance (race + sex) 0.052 (0.150) -0.055 (0.147) -0.123 (0.305) -0.155 (0.093) -0.186 (0.042) -0.176 (0.028) -0.180 (0.023) -0.184 (0.016)
Table 8: Full set of simulation results for the COMPAS dataset for $\eta=0.15$ and for various values of the privacy parameter $\varepsilon$. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 successful runs were performed in each case, except for $\varepsilon=10^{-2}$, which had 23. The final column represents a run without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None
KS Stat 0.540 (0.162) 0.291 (0.083) 0.104 (0.033) 0.047 (0.016) 0.021 (0.005) 0.014 (0.005) 0.014 (0.004) 0.008 (0.003)
KS p-val 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.018 (0.039) 0.408 (0.259) 0.783 (0.248) 0.785 (0.243) 0.974 (0.068)
Cumulative 1-way TVD 1.310 (0.431) 0.579 (0.165) 0.189 (0.046) 0.059 (0.020) 0.021 (0.005) 0.012 (0.004) 0.009 (0.002) 0.008 (0.003)
Cumulative 2-way TVD 5.036 (1.437) 2.462 (0.692) 1.071 (0.168) 0.537 (0.060) 0.217 (0.025) 0.135 (0.013) 0.108 (0.008) 0.100 (0.010)
Cumulative 3-way TVD 8.506 (2.026) 4.511 (1.214) 2.298 (0.285) 1.289 (0.149) 0.572 (0.057) 0.427 (0.029) 0.387 (0.018) 0.301 (0.023)
COD (race) 0.011 (0.036) -0.011 (0.032) -0.015 (0.039) -0.050 (0.029) -0.062 (0.011) -0.066 (0.006) -0.066 (0.005) -0.065 (0.006)
COD (sex) -0.000 (0.024) -0.013 (0.036) -0.003 (0.044) -0.024 (0.031) -0.045 (0.015) -0.047 (0.009) -0.047 (0.009) -0.049 (0.010)
COD (race + sex) 0.010 (0.031) -0.020 (0.105) -0.013 (0.041) -0.039 (0.032) -0.046 (0.017) -0.041 (0.016) -0.040 (0.018) -0.044 (0.013)
Accuracy 0.518 (0.050) 0.521 (0.064) 0.576 (0.060) 0.653 (0.020) 0.670 (0.007) 0.672 (0.005) 0.672 (0.005) 0.672 (0.005)
F1-score 0.483 (0.275) 0.506 (0.248) 0.624 (0.087) 0.694 (0.021) 0.704 (0.011) 0.707 (0.008) 0.707 (0.007) 0.707 (0.007)
TP rate 0.577 (0.393) 0.590 (0.366) 0.660 (0.168) 0.711 (0.054) 0.708 (0.022) 0.713 (0.016) 0.713 (0.015) 0.714 (0.014)
TN rate 0.444 (0.401) 0.434 (0.366) 0.472 (0.228) 0.581 (0.080) 0.623 (0.019) 0.621 (0.009) 0.621 (0.008) 0.620 (0.007)
FN rate 0.423 (0.393) 0.410 (0.366) 0.340 (0.168) 0.289 (0.054) 0.292 (0.022) 0.287 (0.016) 0.287 (0.015) 0.286 (0.014)
FP rate 0.556 (0.401) 0.566 (0.366) 0.528 (0.228) 0.419 (0.080) 0.377 (0.019) 0.379 (0.009) 0.379 (0.008) 0.380 (0.007)
AUC 0.507 (0.092) 0.522 (0.100) 0.589 (0.089) 0.690 (0.026) 0.711 (0.002) 0.713 (0.001) 0.713 (0.001) 0.713 (0.002)
SPD (race) 0.000 (0.153) -0.042 (0.116) -0.122 (0.306) -0.221 (0.089) -0.180 (0.049) -0.191 (0.028) -0.195 (0.029) -0.191 (0.028)
SPD (sex) -0.071 (0.191) -0.046 (0.163) -0.096 (0.218) -0.126 (0.096) -0.169 (0.033) -0.171 (0.026) -0.178 (0.023) -0.178 (0.021)
SPD (race + sex) -0.069 (0.186) -0.053 (0.153) -0.146 (0.265) -0.163 (0.103) -0.174 (0.051) -0.174 (0.047) -0.187 (0.032) -0.185 (0.036)
AOD (race) 0.002 (0.152) -0.037 (0.117) -0.104 (0.304) -0.183 (0.089) -0.139 (0.048) -0.149 (0.026) -0.154 (0.027) -0.150 (0.026)
AOD (sex) -0.067 (0.193) -0.043 (0.168) -0.078 (0.221) -0.092 (0.098) -0.133 (0.032) -0.135 (0.026) -0.142 (0.024) -0.143 (0.021)
AOD (race + sex) -0.067 (0.186) -0.047 (0.161) -0.124 (0.274) -0.119 (0.110) -0.133 (0.050) -0.134 (0.049) -0.147 (0.032) -0.145 (0.036)
FN rate balance (race) -0.010 (0.146) 0.045 (0.112) 0.110 (0.295) 0.183 (0.090) 0.121 (0.053) 0.130 (0.032) 0.135 (0.034) 0.130 (0.034)
FN rate balance (sex) 0.077 (0.182) 0.046 (0.147) 0.090 (0.211) 0.097 (0.094) 0.127 (0.036) 0.128 (0.026) 0.135 (0.023) 0.135 (0.023)
FN rate balance (race + sex) 0.066 (0.176) 0.058 (0.133) 0.154 (0.236) 0.163 (0.093) 0.149 (0.057) 0.148 (0.045) 0.160 (0.034) 0.158 (0.038)
FP rate balance (race) -0.006 (0.162) -0.030 (0.130) -0.097 (0.317) -0.184 (0.093) -0.156 (0.045) -0.169 (0.022) -0.173 (0.021) -0.170 (0.020)
FP rate balance (sex) -0.058 (0.206) -0.041 (0.190) -0.066 (0.238) -0.086 (0.106) -0.139 (0.032) -0.142 (0.028) -0.150 (0.026) -0.150 (0.021)
FP rate balance (race + sex) -0.069 (0.203) -0.037 (0.203) -0.094 (0.321) -0.074 (0.141) -0.117 (0.053) -0.119 (0.056) -0.134 (0.036) -0.132 (0.039)
Table 9: Full set of simulation results for the COMPAS dataset for $\eta=0.08$ and for various values of the privacy parameter $\varepsilon$. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (standard deviation)” of the output values from many runs of the evaluations. At least 30 successful runs were performed in each case, except for $\varepsilon=10^{-2}$ and $\varepsilon=10^{-1.5}$, which had 25 and 26 successful runs, respectively. The final column represents a run without the DP synthesis step as a baseline.